OpenStudy (anonymous):

Is the proportion X/8=5/7 ? Identify the proportion listed below that solves this problem. It takes 5 men 7 days to put up the walls of a house. How many days would it have taken for 8 men to put up the walls of the same house?

OpenStudy (lalaly):

5men 7days 8men x days 5x=7*8\[\frac{x}{8}=\frac{7}{5}\]

OpenStudy (anonymous):

I got x/8=5/7

OpenStudy (lalaly):

no its 7/5